# Graph-Structured Driven Dual Adaptation for Mitigating Popularity Bias

### v1

@@ Abstract
To address these issues, we propose  Graph-Structured Dual Adaptation Framework (GSDA). First, our theoretical analysis reveals a critical limitation of supervised alignment strategies arising from the over-smoothing effect in GCNs. Specifically, as GCNs layers deepen, the distinctiveness between popular and unpopular items gradually diminishes, which can be quantitatively measured by the reduction in their conditional entropy. This progressive loss of distinctiveness severely weakens the effectiveness of supervised alignment methods in alleviating popularity bias at deeper layers. Motivated by this theoretical insight,  GSDA  simultaneously mines structural characteristics and distribution information from the graph adjacency matrix and incorporates a dual adaptive strategy. First, the hierarchical adaptive alignment mechanism uses the F-norm of the adjacency matrix to perform layer-specific weight decay, mitigating the negative effects of deep conditional entropy decay. Second, the distribution-aware dynamic contrast weighting strategy, based on a real-time Gini coefficient, removes the dependence on fixed hyperparameters and enables the model to adapt to diverse data distributions. Extensive experiments on three benchmark datasets show that  GSDA  not only significantly alleviates popularity bias but also consistently outperforms state-of-the-art methods in recommendation performance.
Index Terms:
Recommender Systems, Popularity Bias, Supervised Alignment, Re-weighting, Contrastive Learning
@@ II-A Recommendation Task
Recommender systems aim to predict user preferences toward unobserved items based on historical user-item interactions. Collaborative Filtering (CF), a widely adopted paradigm, predicts these interactions by aligning user and item representations. User-item interactions are represented by a binary matrix $\mathcal{R}\in\{0,1\}^{M\times N}$, where $\mathcal{R}_{u,i}=1$ indicates an interaction between user $u$ and item $i$, and 00 otherwise. Let $U$ and $I$ denote user and item sets with $|U|=M$ and $|I|=N$, respectively. CF models typically learn user embeddings $\mathbf{Z}\in\mathbb{R}^{M\times D}$ and item embeddings $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. The predicted preference score for user $u$ towards item $i$ is calculated as $s(u,i)=\mathbf{z}_{u}^{\top}\mathbf{e}_{i}$, where $\mathbf{z}_{u}$ and $\mathbf{e}_{i}$ are the embeddings for user $u$ and item $i$.
@@ II-B Graph Convolutional Networks
Graph Convolutional Networks (GCNs) have been widely adopted in collaborative filtering due to their capability of modeling high-order collaborative signals between users and items through graph-based information propagation [23, 24, 25]. Specifically, user-item interactions can be naturally represented as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathbf{A})$, where nodes include users and items, and edges represent interactions. The corresponding adjacency matrix $\mathbf{A}\in\mathbb{R}^{(M+N)\times(M+N)}$ is defined as:
@@ III-A Theoretical Analyses
$H(\mathbf{X}_{up}|\mathbf{X}_{p})=-\mathbb{E}_{P(\mathbf{X}_{up},\mathbf{X}_{p})}[\log P(\mathbf{X}_{up}|\mathbf{X}_{p})],$ (8)
where $\mathbf{X}_{up}$ and $\mathbf{X}_{p}$ represent embeddings of unpopular and popular items, respectively. This entropy-based metric directly reflects alignment efficacy—higher values indicate more effective information transfer. In particular, a higher conditional entropy indicates greater uncertainty about unpopular item embeddings, thus implying more potential for beneficial information transfer from popular to unpopular items. Conversely, lower conditional entropy implies that embeddings of unpopular items already possess similar information to popular items, limiting alignment benefits.
@@
This convergence indicates that propagation weights between any node pairs become highly similar as layers deepen. Combining this nearly uniform propagation structure with initially random Gaussian embeddings $\mathbf{X}^{(0)}$ results in the homogenization of node embeddings. Such embedding homogenization, known as over-smoothing, significantly reduces the discriminative power of node representations, especially impacting unpopular items with sparse interactions.
Layer-Wise Conditional Entropy Reduction. Given the homogenization of embeddings induced by over-smoothing, we now analyze how this phenomenon directly leads to a monotonic reduction in conditional entropy. By explicitly deriving the relationship between adjacency matrix differences and entropy, we demonstrate that alignment effectiveness inevitably weakens as the number of layers increases. We now explicitly connect the over-smoothing phenomenon described above to the progressive reduction in conditional entropy observed in supervised alignment. Consider the embeddings of an unpopular item $\mathbf{x}_{up}^{(l)}$ and a popular item $\mathbf{x}_{p}^{(l)}$ at layer $l$, which can be expressed based on GCN propagation:
$\mathbf{x}_{up}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{up,j}\mathbf{x}_{j}^{(0)},\quad\mathbf{x}_{p}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{p,j}\mathbf{x}_{j}^{(0)},$ (10)
@@
To quantify how well popular item embeddings can inform unpopular item embeddings, we use conditional entropy as defined Eq. (8). For tractable analysis, we approximate the conditional probability $P(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})$ via Gaussian kernel density estimation (KDE) [30]:
$P(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})\;\propto\;\exp\Bigl{(}-\frac{\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}}{2\sigma^{2}}\Bigr{)},$ (11)
where $\sigma$ is a bandwidth parameter controlling the smoothness of the probability estimation. By substituting Eq. (11) into the conditional entropy definition, we explicitly establish a relationship between entropy and embedding distance:
$H(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})\;\propto\;\mathbb{E}\bigl{[}\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}\bigr{]}.$ (12)
@@
$\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}\;=\;\sum_{j,k}\,\bigl{(}\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr{)}^{2}\,\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}.$ (13)
@@
$H\bigl{(}\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)}\bigr{)}\;\propto\;\sum_{j,k}\bigl{(}\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr{)}^{2}\mathbb{E}\bigl{[}\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}\bigr{]}.$ (14)
Crucially, as layer depth $l$ increases, the phenomenon of over-smoothing causes $\hat{\mathbf{A}}^{l}_{up,j}$ and $\hat{\mathbf{A}}^{l}_{p,k}$ to become increasingly similar for all node pairs (Eq. (9)). Consequently, as these adjacency matrix differences shrink, the squared embedding distances also decrease, directly causing a layer-wise monotonic reduction in conditional entropy:
$H\bigl{(}\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)}\bigr{)}\;\geq\;H\bigl{(}\mathbf{x}_{up}^{(l+1)}\mid\mathbf{x}_{p}^{(l+1)}\bigr{)}.$ (15)
From this monotonic trend, we conclude that repeated propagation via higher-order adjacency matrices leads to embedding homogenization, which progressively reduces conditional entropy and thus weakens the effectiveness of supervised alignment at deeper GCN layers. Our theoretical findings not only explain why deeper GCN layers diminish alignment effectiveness but also highlight the importance of designing adaptive strategies that dynamically adjust alignment contributions across layers. This theoretical insight serves as a foundation for developing novel approaches to mitigate the impact of embedding homogenization in graph-based learning.
@@ III-B Experimental Validation
We perform experiments on the Gowalla dataset to empirically examine the relationship among over-smoothing, embedding homogenization, and supervised alignment effectiveness in deeper layers. First, we randomly sample 1,000 pairs of popular and unpopular items and calculate their conditional entropy ($H(\mathbf{X}_{up}^{(l)}|\mathbf{X}_{p}^{(l)})$) and cosine similarity across different layers of LightGCN. As shown in Fig.2a, the conditional entropy between popular and unpopular items significantly decreases as the layer depth increases, coinciding with increased embedding similarity. This observation empirically confirms our theoretical finding that *embedding homogenization occurs due to the progressive uniformity of the adjacency matrix entries in deeper layers (Eq. (9)), which directly reduces the distinguishability of unpopular item embeddings.
@@ IV-A Adaptive Hierarchical Supervised Alignment
$\mathcal{L}_{sa}=\frac{1}{L+1}\sum_{l=0}^{L}w_{l}\left(\frac{1}{|U|}\sum_{u\in\mathcal{U}}\frac{1}{|\boldsymbol{I}_{u}|}\sum_{i\in\boldsymbol{I}_{u}^{up},j\in\boldsymbol{I}_{u}^{p}}\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}\right),$ (19)
where $U$ denotes the set of users, and $\boldsymbol{I}_{u}^{p}$ and $\boldsymbol{I}_{u}^{up}$ are the sets of popular and unpopular items interacted with by user $u$, respectively. The term $\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}$ explicitly aligns embeddings of popular and unpopular items at each layer, with weights adaptively adjusted to reflect the structural significance captured at different depths.
@@ IV-B Adaptive Re-weighting Contrastive Learning
Specifically, rather than assigning uniform or fixed weights, we introduce a user-specific weighting strategy that more accurately captures each user’s contribution to item popularity. Each user $u$ is assigned a weight inversely proportional to their degree (number of interactions), formally defined as:
$w_{u}=\frac{1}{d_{u}+\varepsilon},$ (20)
where $d_{u}$ is the degree of user $u$, and $\varepsilon$ is a small constant introduced to maintain numerical stability, preventing division by zero when users have minimal interactions. By introducing this user-specific weighting, the contributions from highly active users are balanced, ensuring a more accurate and nuanced estimation of item popularity.
@@
$\mathcal{L}_{cl}=(1-w)\mathcal{L}_{cl}^{p}+w\mathcal{L}_{cl}^{up},$ (23)
where $\mathcal{L}_{cl}^{p}$ and $\mathcal{L}_{cl}^{up}$ represent the contrastive losses computed separately for popular and unpopular items. By adaptively adjusting weights based on $w$, our method dynamically reduces the dominance of the over-represented group. When the distribution is highly skewed (large $w$), the model strategically reduces the influence of popular items to enhance embedding diversity and robustness. Conversely, when item popularity is relatively balanced (small $w$), the method treats both groups more evenly, maintaining optimal performance.
@@ IV-C Model Optimization and Algorithm
where $\mathcal{L}_{rec}$ is the base recommendation loss, capturing fundamental user–item preference signals (e.g., BPR [14] or another ranking objective); $\mathcal{L}_{sa}$ is the adaptive hierarchical supervised alignment loss, designed to address the over-smoothing effect in deeper GCN layers and preserve discriminative item representations;$\mathcal{L}_{cl}$ is the adaptive re-weighting contrastive loss, derived from dynamic item popularity estimation, enabling the model to flexibly adjust sample weights based on real-time popularity distributions. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of $\mathcal{L}_{sa}$ and $\mathcal{L}_{cl}$, respectively. After training, we perform inference by computing the dot product between the final user and item embeddings, ranking items based on their predicted relevance [26]. The detailed algorithm for  GSDA  is presented in Algorithm 1.
@@
19:        Compute $\mathcal{L}_{cl}\leftarrow(1-w)\mathcal{L}_{cl}^{p}+w\mathcal{L}_{cl}^{up}$ (Eq. 23 )

### v2

@@ Abstract
To address these challenges, we propose  Graph-Structured Dual Adaptation Framework (GSDA), a dual adaptive framework for mitigating popularity bias in recommendation. Our theoretical analysis shows that supervised alignment in GCNs is hindered by the over-smoothing effect, where the distinction between popular and unpopular items diminishes as layers deepen, reducing the effectiveness of alignment at deeper levels. To overcome this limitation,  GSDAintegrates a hierarchical adaptive alignment mechanism that counteracts entropy decay across layers together with a distribution-aware contrastive weighting strategy based on the Gini coefficient, enabling the model to adapt its debiasing strength dynamically without relying on fixed hyperparameters. Extensive experiments on three benchmark datasets demonstrate that  GSDAeffectively alleviates popularity bias while consistently outperforming state-of-the-art methods in recommendation performance. The source code for our method is available at https://github.com/miaomiao-cai2/GSDA.
@@ II-A Recommendation Task
Collaborative Filtering (CF), a widely adopted paradigm, predicts these interactions by aligning user and item representations. User-item interactions are represented by a binary matrix $\mathcal{R}\in\{0,1\}^{M\times N}$, where $\mathcal{R}_{u,i}=1$ indicates an interaction between user $u$ and item $i$, and 0 otherwise. Let $U$ and $I$ denote user and item sets with $|U|=M$ and $|I|=N$, respectively. CF models typically learn user embeddings $\mathbf{Z}\in\mathbb{R}^{M\times D}$ and item embeddings $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. The predicted preference score for user $u$ towards item $i$ is calculated as $s(u,i)=\mathbf{z}_{u}^{\top}\mathbf{e}_{i}$, where $\mathbf{z}_{u}$ and $\mathbf{e}_{i}$ are the embeddings for user $u$ and item $i$.
@@ II-B Graph Convolutional Networks
Graph Convolutional Networks (GCNs) have been widely adopted in collaborative filtering [23, 24, 25]. Specifically, user-item interactions can be naturally represented as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathbf{A})$, where nodes include users and items, and edges represent interactions. The corresponding adjacency matrix $\mathbf{A}\in\mathbb{R}^{(M+N)\times(M+N)}$ is defined as:
@@ III-A Theoretical Analyses
$H(\mathbf{X}_{\text{up}}|\mathbf{X}_{\text{p}})=-\mathbb{E}_{P(\mathbf{X}_{\text{up}},\mathbf{X}_{\text{p}})}[\log P(\mathbf{X}_{\text{up}}|\mathbf{X}_{\text{p}})],$ (8)
where $\mathbf{X}_{\text{up}}$ and $\mathbf{X}_{\text{p}}$ denote embeddings of unpopular and popular items, respectively. This entropy-based metric directly reflects alignment efficacy—higher values indicate more effective information transfer. In particular, a higher conditional entropy indicates greater uncertainty about unpopular item embeddings, thus implying more potential for beneficial information transfer from popular to unpopular items. Conversely, lower conditional entropy implies that embeddings of unpopular items already possess similar information to popular items, limiting alignment benefits. For analytical tractability, we further assume that initial node embeddings are sampled independently from an isotropic Gaussian distribution with zero mean, which allows expected squared distances across nodes to be reduced to a constant factor. This standard assumption in GCN theory simplifies the derivation without affecting the qualitative conclusions about entropy decay.
@@
This convergence requires the user–item graph to be connected and aperiodic; in practice, LightGCN satisfies this condition by including self-loops. This convergence means that, as layers deepen, the propagation weights between any two nodes become almost indistinguishable. When combined with the random initialization of embeddings, this uniformity drives different nodes to become increasingly similar, a phenomenon widely known as over-smoothing. For unpopular items with sparse connections, this loss of distinctiveness is particularly severe.
Layer-Wise Conditional Entropy Reduction. Given the homogenization of embeddings induced by over-smoothing, we now analyze how this phenomenon directly leads to a monotonic reduction in conditional entropy. By explicitly deriving the relationship between adjacency matrix differences and entropy, we demonstrate that alignment effectiveness inevitably weakens as the number of layers increases. We now explicitly connect the over-smoothing phenomenon described above to the progressive reduction in conditional entropy observed in supervised alignment. Consider the embeddings of an unpopular item $\mathbf{X}_{\text{up}}^{(l)}$ and a popular item $\mathbf{X}_{\text{p}}^{(l)}$ at layer $l$, which can be expressed based on GCN propagation:
$\mathbf{X}_{\text{up}}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{up,j}\mathbf{x}_{j}^{(0)},\quad\mathbf{X}_{\text{p}}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{p,j}\mathbf{x}_{j}^{(0)},$ (10)
@@
To quantify how well popular item embeddings can inform unpopular item embeddings, we use conditional entropy as defined Eq. (8). For tractable analysis, we approximate the conditional probability $P(\mathbf{X}_{\text{up}}^{(l)}\mid\mathbf{X}_{\text{p}}^{(l)})$ via Gaussian kernel density estimation (KDE) [30]:
$P(\mathbf{X}_{\text{up}}^{(l)}\mid\mathbf{X}_{\text{p}}^{(l)})\;\propto\;\exp\Bigl(-\frac{\|\mathbf{X}_{\text{up}}^{(l)}-\mathbf{X}_{\text{p}}^{(l)}\|^{2}}{2\sigma^{2}}\Bigr),$ (11)
where $\sigma$ is a kernel bandwidth parameter controlling the smoothness of the density estimation. This KDE surrogate is used solely to relate conditional entropy to squared Euclidean distance. The constant terms introduced by this approximation are independent of the layer index $l$ and therefore do not affect the monotonic trend we establish. By substituting Eq. (11) into the conditional entropy definition, we explicitly establish a relationship between entropy and embedding distance:
$H(\mathbf{X}_{\text{up}}^{(l)}\mid\mathbf{X}_{\text{p}}^{(l)})\;\propto\;\mathbb{E}\bigl[\|\mathbf{X}_{\text{up}}^{(l)}-\mathbf{X}_{\text{p}}^{(l)}\|^{2}\bigr].$ (12)
@@
$\|\mathbf{X}_{\text{up}}^{(l)}-\mathbf{X}_{\text{p}}^{(l)}\|^{2}\;=\;\sum_{j,k}\,\bigl(\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr)^{2}\,\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}.$ (13)
@@
$\small H\bigl(\mathbf{X}_{\text{up}}^{(l)}\mid\mathbf{X}_{\text{p}}^{(l)}\bigr)\;\propto\;\sum_{j,k}\bigl(\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr)^{2}\mathbb{E}\bigl[\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}\bigr].$ (14)
This entropy decreases monotonically with layer depth $l$, reflecting the diminishing distinctiveness of unpopular items. As the number of layers increases, over-smoothing makes adjacency entries between popular and unpopular items progressively similar. This shrinking difference reduces embedding distances, which in turn leads to a monotonic decline in conditional entropy across layers. Put simply, deeper propagation erodes the unique information of unpopular items, weakening the benefit of alignment. Unlike prior over-smoothing analyses [31, 32, 33], which treat all nodes symmetrically, our analysis explicitly incorporates the degree heterogeneity of user–item bipartite graphs. We show that high-degree (popular) items dominate the leading spectral components of $\hat{\mathbf{A}}$, causing low-degree (long-tail) items to collapse towards them. This provides a new perspective: over-smoothing is not only a homogenization phenomenon but also a mechanism that amplifies popularity bias.
$H\bigl(\mathbf{X}_{\text{up}}^{(l)}\mid\mathbf{x}_{p}^{(l)}\bigr)\;\geq\;H\bigl(\mathbf{X}_{\text{up}}^{(l+1)}\mid\mathbf{x}_{p}^{(l+1)}\bigr).$ (15)
In summary, repeated propagation through higher-order adjacency matrices homogenizes embeddings, progressively reducing conditional entropy and weakening the effectiveness of supervised alignment in deeper GCN layers.
@@ III-B Experimental Validation
We perform experiments on the Gowalla dataset to empirically examine the relationship among over-smoothing, embedding homogenization, and supervised alignment effectiveness in deeper layers. First, we randomly sample 1,000 pairs of popular and unpopular items and calculate their conditional entropy ($H(\mathbf{X}_{\text{up}}^{(l)}|\mathbf{X}_{p}^{(l)})$) and cosine similarity across different layers of LightGCN. As shown in Fig.2a, the conditional entropy between popular and unpopular items significantly decreases as the layer depth increases, coinciding with increased embedding similarity. This observation empirically confirms our theoretical finding that *embedding homogenization occurs due to the progressive uniformity of the adjacency matrix entries in deeper layers (Eq. (9)), which directly reduces the distinguishability of unpopular item embeddings.
@@ IV-A Adaptive Hierarchical Supervised Alignment
$\small\mathcal{L}_{sa}=\frac{1}{L+1}\sum_{l=0}^{L}w_{l}\left(\frac{1}{|U|}\sum_{u\in\mathcal{U}}\frac{1}{|\boldsymbol{I}_{u}|}\sum_{i\in\boldsymbol{I}_{u,\text{up}},j\in\boldsymbol{I}_{u,\text{p}}}\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}\right),$ (19)
where $U$ denotes the set of users, and $\boldsymbol{I}_{u,\text{p}}$ and $\boldsymbol{I}_{u,\text{up}}$ are the sets of popular and unpopular items interacted with by user $u$, respectively. The term $\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}$ explicitly aligns embeddings of popular and unpopular items at each layer, with weights adaptively adjusted to reflect the structural significance captured at different depths.
@@ IV-B Adaptive Re-weighting Contrastive Learning
Specifically, rather than assigning uniform or fixed weights, we introduce a user-specific weighting strategy that more accurately captures each user’s contribution to item popularity. Each user $u$ is assigned a weight inversely proportional to the logarithm of their degree (number of interactions), formally defined as:
$w_{u}=\frac{1}{\log(1+d_{u})+\varepsilon},$ (20)
where $d_{u}$ is the degree of user $u$, and $\varepsilon$ is a small constant introduced to maintain numerical stability. This log-based formulation provides a smoother penalty for highly active users compared to the linear inverse design, thereby avoiding excessive down-weighting. By adopting this user-specific weighting, the influence of frequent users is moderated in a more balanced manner, leading to a more stable and nuanced estimation of item popularity.
@@
$\mathcal{L}_{cl}=(1-w)\mathcal{L}_{cl}^{\text{p}}+w\mathcal{L}_{cl}^{\text{up}},$ (23)
where $\mathcal{L}_{cl}^{\text{p}}$ and $\mathcal{L}_{cl}^{\text{up}}$ represent the contrastive losses computed separately for popular and unpopular items. By adaptively adjusting weights based on $w$, our method dynamically reduces the dominance of the over-represented group. When the distribution is highly skewed (large $w$), the model strategically reduces the influence of popular items to enhance embedding diversity and robustness. Conversely, when item popularity is relatively balanced (small $w$), the method treats both groups more evenly, maintaining optimal performance.
@@ IV-C Model Optimization and Algorithm
[…]
@@
19:   Compute $\mathcal{L}_{cl}\leftarrow(1-w)\mathcal{L}_{cl}^{p}+w\mathcal{L}_{cl}^{\text{up}}$ (Eq. 23 )
@@
IV-D Model Analysis
IV-D1 Space Complexity
GSDA follows LightGCN in maintaining user and item embeddings, requiring $(M+N)d$ parameters, where $M$ and $N$ are the numbers of users and items, and $d$ is the embedding dimension. During training, the adaptive hierarchical supervised alignment module stores intermediate embeddings for $L$ layers, incurring $\mathcal{O}(L(M+N)d)$ additional memory. Since $L$ is typically small, this overhead is modest. The adaptive re-weighting contrastive learning module only stores lightweight popularity statistics (e.g., user degrees, item popularity scores, Gini coefficient), which are negligible compared to embedding storage.
IV-D2 Time Complexity
The main computational cost of GSDA arises from the sparse embedding propagation in $L$ layers, identical to LightGCN, with complexity $\mathcal{O}(|R^{+}|Ld)$, where $|R^{+}|$ is the number of non-zero entries in the interaction matrix. The adaptive hierarchical supervised alignment adds $\mathcal{O}(L|\mathcal{B}|d)$ cost for aligning popular–unpopular item pairs in a minibatch $\mathcal{B}$; in practice, a small constant number of pairs per user is sampled, so this remains linear in the batch size. The adaptive re-weighting contrastive learning introduces a similar $\mathcal{O}(L|\mathcal{B}|d)$ term for similarity computation. Popularity statistics and the Gini coefficient are computed once per epoch with negligible cost. Therefore, the overall per-epoch complexity is $\mathcal{O}(|R^{+}|Ld+L|\mathcal{B}|d)$, which is comparable to existing GCN-based debiasing methods and practical in real-world settings.
IV-D3 Limitations
Although GSDA achieves consistent improvements in debiasing and recommendation accuracy, its performance can be influenced by certain data characteristics. The adaptive hierarchical alignment assumes that the user–item graph provides sufficiently informative structural signals; in extremely sparse or highly irregular graphs, these signals may be weak, reducing alignment effectiveness. The adaptive re-weighting contrastive learning relies on accurate popularity estimation, and in highly dynamic environments with abrupt popularity shifts, frequent re-estimation may introduce transient instability and extra computation. Moreover, GSDA is designed for bipartite collaborative filtering graphs without side information; in cold-start or heterogeneous scenarios, incorporating additional modalities or graph types may be necessary to maintain robustness. Future work could explore temporal smoothing of popularity estimation, efficient incremental graph updates, and extensions to multi-relational or multimodal settings.